# Shadows aspect of rotating black holes in the Einstein-AdS SU(N)-nonlinear sigma model

### v1

@@ III.2 The BH shadow
Escaping photons in the unstable orbit give the possibility for distant observers to backward-trace and obtain a shadow cast for an observer located at the spatial position $(r_{o},\theta_{o})$. Such an observer is also known as the Zero Angular Momentum Observers (ZAMOs), for whom the shadow is seen in the celestial plane, identified by the coordinates Johannsen:2013vgc
@@
It is evident that a decrease in the inclination angle $\theta_{o}$ leads to an increase in the size of the BH shadow. Additionally, as expected, an increase in the flavor number $N$ also results in the enlargement of the shadow, as demonstrated in the diagrams (a)–(d) of Fig. 9.
@@ IV Shadow observables and BH parameter estimation
To have BH models that are observationally consistent, the parameters associated with theoretical BHs need to be constrained using EHT observations. The observed images of the supermassive BH M87* and Sgr A* are consistent with the predictions of the Kerr BH of general relativity. In the following, we identify key observables and examine their behavior within the framework of the RASN-BH.
@@ IV.1 Distortion
In fact, the distortion measures the deviation of the BH shadow’s shape from a circular reference. In other words, it quantifies the extent to which the shadow of a rotating BH departs from circularity. Here, we adopt the method proposed in Ref. Hioki:2009na , where the linear radius of the shadow is defined as
@@ IV.2 Parameter estimation
The shadow of a BH reveals key characteristics of the underlying spacetime, such as its shape and size. As such, it provides a valuable tool for testing new gravity theories and for constraining BH parameters (jusufi_constraints_2022, ; okyay_nonlinear_2022, ; afrin_parameter_2021, ; chen_superradiant_2022, ; afrin_tests_2023, ; sarikulov_shadow_2022, ; zahid_shadow_2023, ; zubair_4d_2023, ). Estimations of BH parameters can be derived from the observables associated with the shadow. To achieve this, it is necessary to define the key parameters that describe the shadow’s size and shape. In addition to the previously discussed method, a more direct approach can be employed, utilizing a coordinate-independent formalism (abdujabbarov_coordinate-independent_2015, ; kumar_black_2020, ) and focusing on shadow observables such as the shadow’s area and oblateness. The area of the BH shadow, denoted $A_{s}$, and its oblateness, $D_{s}$, provide insights into the deformation of the shadow and are defined as (kumar_black_2020, )
@@ IV.3 Constraints from the EHT observations for M87* and Sgr A*
The shadow images of two supermassive BHs, M87* and Sgr A*, captured by the EHT collaboration (EventHorizonTelescope:2019dse, ; EventHorizonTelescope:2019ggy, ; EventHorizonTelescope:2022wkp, ; EventHorizonTelescope:2022xqj, ), provide compelling motivation for studying BH shadows, which offer significant scientific insights. These observations enable us to estimate the parameters of BHs within the context of various modified or alternative gravity models, helping to identify which models are most consistent with the observed data. In this section, we apply the constraints derived from the EHT observations of M87* and Sgr A* to analyze the parameters of our rotating BH model. Specifically, we utilize the angular diameters of these BHs, as measured by the EHT collaboration, to constrain the model. The angular diameter of a BH’s shadow, observed from a distance $d$, can be expressed as (afrin_parameter_2021, ; zahid_shadow_2023, )
@@
where $R_{a}$ represents the areal radius of the shadow, and $d$ is the observer’s distance to the BH. Using Eq. (33), we find that the angular diameter of the shadow depends on the BH parameters and the observation angle, with an implicit dependence on the BH mass. In our analysis, we treat the supermassive BHs M87* and Sgr A* as RASN-BHs and compare the theoretical shadow predictions with the observational data from the EHT. The mass of M87* and its distance from Earth are taken as $M=6.5\times 10^{9}M_{\odot}$ and $d=16.8\,\text{Mpc}$, respectively EventHorizonTelescope:2019pgp ; EventHorizonTelescope:2019ggy . For simplicity, the uncertainties associated with the mass and distance measurements of these supermassive BHs are not considered in the calculations. The angular diameter of the shadow image for M87* is reported as $\theta_{d}=(42\pm 3)\,\mu\text{as}$ at the $1\sigma$ confidence level EventHorizonTelescope:2019dse . In Fig. 14, we illustrate the mutual behavior of $a/M$ and $K$ using the theoretical angular diameter in Eq. (35), for the observational data inferred for M87*, considering two distinct values for the $N$-parameter and two inclinations, $\theta_{o}=90^{\circ}$ and $17^{\circ}$. Based on the EHT observational results from the shadow images of M87*, it has been shown that the spin parameter of M87* is approximately $a\approx(0.9\pm 0.05)M$ Tamburini:2019vrf , while $\theta_{o}=17^{\circ}$ Daly:2023axh .
@@
Similarly, constraints for the shadow image of Sgr A* can be derived from EHT observations. The angular diameter of the shadow of the supermassive BH Sgr A* is reported as $\theta_{o}=(48.7\pm 7)\,\mu\text{as}$ EventHorizonTelescope:2022wkp . The mass of Sgr A* and its distance from the solar system are approximately $M=4\times 10^{6}M_{\odot}$ and $d=8\,\text{kpc}$, respectively EventHorizonTelescope:2022exc ; EventHorizonTelescope:2022xqj . In Fig. 15, the constraints obtained from Sgr A* are illustrated for the two cases under consideration. The fitted density plots of the angular diameter of the BH fall within the measured angular diameter of Sgr A*, $\theta_{d}=(48.7\pm 7)\,\mu\text{as}$. Furthermore, recent analysis of Sgr A* suggests that its spin should be $a=(0.9\pm 0.06)M$ (Daly:2023axh, ).
@@ V The energy emission rate
From a quantum mechanical standpoint, near a BH’s event horizon, particles can be both created and annihilated. Those with positive energy can escape the event horizon through tunneling. This mechanism drives BHs to emit radiation, which could ultimately lead to their evaporation. This phenomenon, known as Hawking radiation, arises due to quantum effects, where BHs radiate thermally, progressively losing mass and energy until they eventually disappear Hawking:1974rv . At high energies, Hawking radiation typically emerges within a finite cross-sectional area, denoted by $\sigma_{l}$. For distant observers far from the BH, this cross-section approaches the shadow cast by the BH belhaj_deflection_2020 ; wei_observing_2013 . It has been established that $\sigma_{l}$ is directly related to the area of the photon ring and can be approximated as wei_observing_2013 ; decanini_fine_2011 ; li_shadow_2020
@@
As seen in the diagrams, faster-rotating BHs emit less energy. Furthermore, for all values of $N$, a larger $K$-parameter results in a reduced energy emission rate. For slower-rotating BHs, the Kerr-AdS BH exhibits the highest evaporation rate. However, this behavior evolves as the spin parameter increases (see diagrams (e) and (f)). Specifically, the energy emission rate profile of the RASN-BH for each value of the $K$-parameter approaches that of the Kerr-AdS BH and eventually surpasses it with an increase in the spin parameter. Additionally, it is evident that as $a$ increases, the profiles become more similar, regardless of the value of $N$.
@@ Appendix A Notes on the derivation of the rotating spacetime
To derive the rotating spacetime (8), we first apply the MNJA to construct the core components of the function (9a), and then incorporate the terms associated with the cosmological constant. The technique follows the approach outlined in Refs. Azreg:2014 ; azreg-ainou_static_2014 .
@@
The MNJA provides an effective means of generating rotating solutions from static ones while ensuring the separability of the Hamilton-Jacobi equation for null geodesics junior_spinning_2020 . This method has been extensively used in various gravity theories Shaikh:2020 ; kumar_rotating_2018 ; contreras_black_2019 ; fathi_spherical_2023 .
Now based on the static spacetime form given in Eqs. (3) and (4), one can infer $\Delta_{r}=a^{2}-2Mr+r^{2}\left[1-\frac{1}{6}\kappa KN\left(N^{2}-1\right)\right]-\frac{1}{3}{\Lambda r^{4}}$. However, to fully incorporate the rotating black hole spacetime, including a cosmological constant, we follow the method outlined in Ref. (xu_kerr-newman-ads_2017, ). Assuming the general stationary line element (8), we explicitly derive the components of the field equations for the Einstein-SU(N)-NLSM within this spacetime. This can be efficiently performed using the computation packages such as RGTC (RGTC, ) in Mathematica®. By carrying out this procedure, one can verify that the line element (8), together with the expressions in Eqs. (9), satisfies the field equations of the theory. Consequently, we conclude that the RASN-BH represents a stationary solution to the model.

### v2

@@ III.2 The BH shadow
Escaping photons in the unstable orbit give the possibility for distant observers to backward-trace and obtain a shadow cast for an observer located at the spatial position $(r_{o},\theta_{o})$. Such an observer is also known as the Zero Angular Momentum Observers (ZAMOs), for whom the shadow is seen in the celestial plane, identified by the coordinates [156]
@@
It is evident that a decrease in the inclination angle $\theta_{o}$ leads to an increase in the size of the BH shadow. Additionally, as expected, an increase in the flavor number $N$ also results in the enlargement of the shadow, as demonstrated in the diagrams (a)–(d) of Fig. 9. It is worth mentioning that a larger shadow may indicate a weaker gravitational field near the black hole’s horizon, potentially suggesting gravitational stability. However, without more detailed studies on the black hole’s response to perturbations, it is not possible to make definitive statements regarding its stability.
@@ IV Shadow observables and BH parameter estimation
To have BH models that are observationally consistent, the parameters associated with theoretical BHs need to be constrained using EHT observations. The observed images of the supermassive BH M87* and Sgr A* are consistent with the predictions of the Kerr BH of general relativity. In the following, we identify key observables and examine their behavior within the framework of the RASN-BH. Note that several studies have explored constraints on the parameters of general black holes derived from the NJA (see, for example, Refs. [157, 158]). In this section, we focus on the most salient aspects of the black hole, whose constraints provide valuable insights into its fundamental characteristics.
@@ IV.1 Distortion
In fact, the distortion measures the deviation of the BH shadow’s shape from a circular reference. In other words, it quantifies the extent to which the shadow of a rotating BH departs from circularity. Here, we adopt the method proposed in Ref. [48], where the linear radius of the shadow is defined as
@@ IV.2 Parameter estimation
The shadow of a BH reveals key characteristics of the underlying spacetime, such as its shape and size. As such, it provides a valuable tool for testing new gravity theories and for constraining BH parameters [159, 160, 161, 162, 163, 164, 165, 166]. Estimations of BH parameters can be derived from the observables associated with the shadow. To achieve this, it is necessary to define the key parameters that describe the shadow’s size and shape. In addition to the previously discussed method, a more direct approach can be employed, utilizing a coordinate-independent formalism [167, 168] and focusing on shadow observables such as the shadow’s area and oblateness. The area of the BH shadow, denoted $A_{s}$, and its oblateness, $D_{s}$, provide insights into the deformation of the shadow and are defined as [168]
@@ IV.3 Constraints from the EHT observations for M87* and Sgr A*
The shadow images of two supermassive BHs, M87* and Sgr A*, captured by the EHT collaboration [17, 19, 27, 28], provide compelling motivation for studying BH shadows, which offer significant scientific insights. These observations enable us to estimate the parameters of BHs within the context of various modified or alternative gravity models, helping to identify which models are most consistent with the observed data. In this section, we apply the constraints derived from the EHT observations of M87* and Sgr A* to analyze the parameters of our rotating BH model. Specifically, we utilize the angular diameters of these BHs, as measured by the EHT collaboration, to constrain the model. The angular diameter of a BH’s shadow, observed from a distance $d$, can be expressed as [161, 165]
@@
where $R_{a}$ represents the areal radius of the shadow, and $d$ is the observer’s distance to the BH. Using Eq. (33), we find that the angular diameter of the shadow depends on the BH parameters and the observation angle, with an implicit dependence on the BH mass. In our analysis, we treat the supermassive BHs M87* and Sgr A* as RASN-BHs and compare the theoretical shadow predictions with the observational data from the EHT. The mass of M87* and its distance from Earth are taken as $M=6.5\times 10^{9}M_{\odot}$ and $d=16.8\,\text{Mpc}$, respectively [18, 19]. For simplicity, the uncertainties associated with the mass and distance measurements of these supermassive BHs are not considered in the calculations. The angular diameter of the shadow image for M87* is reported as $\theta_{d}=(42\pm 3)\,\mu\text{as}$ at the $1\sigma$ confidence level [17]. In Fig. 14, we illustrate the mutual behavior of $a/M$ and $K$ using the theoretical angular diameter in Eq. (35), for the observational data inferred for M87*, considering two distinct values for the $N$-parameter and two inclinations, $\theta_{o}=90^{\circ}$ and $17^{\circ}$. Based on the EHT observational results from the shadow images of M87*, it has been shown that the spin parameter of M87* is approximately $a\approx(0.9\pm 0.05)M$ [169], while $\theta_{o}=17^{\circ}$ [170].
@@
Similarly, constraints for the shadow image of Sgr A* can be derived from EHT observations. The angular diameter of the shadow of the supermassive BH Sgr A* is reported as $\theta_{o}=(48.7\pm 7)\,\mu\text{as}$ [27]. The mass of Sgr A* and its distance from the solar system are approximately $M=4\times 10^{6}M_{\odot}$ and $d=8\,\text{kpc}$, respectively [23, 28]. In Fig. 15, the constraints obtained from Sgr A* are illustrated for the two cases under consideration. The fitted density plots of the angular diameter of the BH fall within the measured angular diameter of Sgr A*, $\theta_{d}=(48.7\pm 7)\,\mu\text{as}$. Furthermore, recent analysis of Sgr A* suggests that its spin should be $a=(0.9\pm 0.06)M$ [170].
@@ V The energy emission rate
From a quantum mechanical standpoint, near a BH’s event horizon, particles can be both created and annihilated. Those with positive energy can escape the event horizon through tunneling. This mechanism drives BHs to emit radiation, which could ultimately lead to their evaporation. This phenomenon, known as Hawking radiation, arises due to quantum effects, where BHs radiate thermally, progressively losing mass and energy until they eventually disappear [171]. At high energies, Hawking radiation typically emerges within a finite cross-sectional area, denoted by $\sigma_{l}$. For distant observers far from the BH, this cross-section approaches the shadow cast by the BH [172, 173]. It has been established that $\sigma_{l}$ is directly related to the area of the photon ring and can be approximated as [173, 174, 175]
@@
As seen in the diagrams, faster-rotating BHs emit less energy. Furthermore, for all values of $N$, a larger $K$-parameter results in a reduced energy emission rate. For slower-rotating BHs, the Kerr-AdS BH exhibits the highest evaporation rate. However, this behavior evolves as the spin parameter increases (see diagrams (e) and (f)). Specifically, the energy emission rate profile of the RASN-BH for each value of the $K$-parameter approaches that of the Kerr-AdS BH and eventually surpasses it with an increase in the spin parameter. Additionally, it is evident that as $a$ increases, the profiles become more similar, regardless of the value of $N$. Note also that, as inferred from the diagrams, for a fixed spin parameter, the emission rate decreases with an increase in the flavor number $N$. This suggests that theories with larger $N$ may result in more stable black holes.
@@ Appendix A Notes on the derivation of the rotating spacetime
To derive the rotating spacetime (8), we first apply the MNJA to construct the core components of the function (9a), and then incorporate the terms associated with the cosmological constant. The technique follows the approach outlined in Refs. [112, 176].
@@
The MNJA provides an effective means of generating rotating solutions from static ones while ensuring the separability of the Hamilton-Jacobi equation for null geodesics [177]. This method has been extensively used in various gravity theories [178, 179, 180, 146].
Now based on the static spacetime form given in Eqs. (3) and (4), one can infer $\Delta_{r}=a^{2}-2Mr+r^{2}\left[1-\frac{1}{6}\kappa KN\left(N^{2}-1\right)\right]-\frac{1}{3}{\Lambda r^{4}}$. However, to fully incorporate the rotating black hole spacetime, including a cosmological constant, we follow the method outlined in Ref. [181]. Assuming the general stationary line element (8), we explicitly derive the components of the field equations for the Einstein-SU(N)-NLSM within this spacetime. This can be efficiently performed using the computation packages such as RGTC [182] in Mathematica®. By carrying out this procedure, one can verify that the line element (8), together with the expressions in Eqs. (9), satisfies the field equations of the theory. Consequently, we conclude that the RASN-BH represents a stationary solution to the model.